In a railway yard goods trains arrive that at the rate of 30 trains per day. Assuming
that the inter-arrival time follows an exponential distribution and the service time
distribution is also exponential with an average of 36 minutes, calculate the
following :
(i) The average number of trains in the queue.
(ii) The probability that the queue size is greater than or equal to 10
1
Expert's answer
2019-03-11T12:44:08-0400

(i) Average number or trains in the queue is

E(m)=(λ2 )/(μ(μ-λ))

λ=30/(60x24)=1/48 trains per minute

E(m)=((1/48)2 )/(1/36(1/36-1/48))=108/48=2.25 or rounghly 2 trains

(ii) The probability that number or trains in it system exceeds 10

P(≥10)=P10 =(λ/μ)10=((1/48)/(1/36))10=(0.75)10 =0.06